OpenStudy (anonymous):

evaluate the definite integral of (x+1)^3/2 from 0 to 1

OpenStudy (anonymous):

\(u=x+1, du = dx, u(0)=1, u(1)=2\) compute \[\int_1^2u^{\frac{3}{2}}du\]

OpenStudy (anonymous):